# Multi-Modal Data-Efficient 3D Scene Understanding for Autonomous Driving

### v2

@@ 5 Experiments
[…]
@@
5.1 Datasets
5.2 Experimental Setups
3D Backbone Configurations. To validate that LaserMix++ is universally applicable to different LiDAR representations, we conduct experiments using a total of five backbones, including FIDNet [36] (range view), PolarNet [37], MinkUNet [52] and Cylinder3D [33] (sparse voxel), and SPVCNN [41] (multi-view fusion). The input resolution of range images is set to 32$\times$1920 for nuScenes [38] and 64$\times$2048 for SemanticKITTI [28] and ScribbleKITTI [7]. The grid cell size of bird’s eye view methods is set to [480, 360, 32]. The voxel size of voxel-based methods is fixed as [240, 180, 20] for all three datasets. The same voxel size is applied to the voxel branch in the multi-view fusion backbone.
Implementation Details. Our LaserMix++ framework is established based on the MMDetection3D codebase [107]. We denote the supervised-only baseline as sup.-only in our experiments. For semi-supervised learning, the labeled/unlabeled data are selected in four ways as shown in Figure 4: 1) random sampling, 2) uniform sampling, 3) sequential sampling, and 4) ST-RFD [11] sampling strategies. Due to the lack of previous works, we also compare consistency regularization [18, 21, 61] and entropy minimization [67] methods from semi-supervised image segmentation. The number of spatial area sets $m$ in multi-modal LaserMix is uniformly sampled from 2 to 6. The loss weights $\lambda_{\text{mix}}$, $\lambda_{\text{mt}}$, $\lambda_{\text{c2l}}$, and $\lambda_{\text{lkg}}$ in Equation 16 are set to 2.0, 250, 1.5, and 1.0, respectively. All experiments are implemented using PyTorch on eight NVIDIA A100 GPUs. All baseline models are trained with a batch size of 2 on each GPU, along with the AdamW optimizer [108], OneCycle learning rate schedule [109], and a learning rate of 0.008. We do not include any type of test-time augmentation or model ensemble during the evaluation. For additional details, kindly refer to the Appendix.
Evaluation Metrics. In this work, we report the intersection-over-union (IoU) for each semantic class and the mean IoU (mIoU) scores across all semantic classes in semi-supervised learning benchmarks. For out-of-distribution robustness assessment, we follow Robo3D [31] protocol to report the mean corruption error (mCE) and mean resilience rate (mRR).
5.3 Comparative Study
Improvements over Baselines. In Tab. II, we compare the proposed LaserMix++ with the LaserMix [10] and sup.-only baselines on nuScenes [38], SemanticKITTI [28], and ScribbleKITTI [7]. Since our approaches can be universally applied to different LiDAR representations, we also set up the benchmark using various backbones from the literature. We observe consistent improvements achieved across all different settings by integrating multi-modal learning for data-efficient 3D scene understanding. On average, there is a 2% to 3% performance gain brought by LaserMix++ over the previous framework. The improvements are especially predominant when the number of labeled data is extremely limited (e.g., 1%), which verifies the effectiveness of our approaches. As for the sup.-only baselines, the gains vary under different backbones. LaserMix++ yields up to 13% mIoU improvements across three datasets when using the range view backbone [36]. Similar trends hold for the bird’s eye view [37], sparse voxel [33, 52], and multi-view fusion [41] backbones, where the mIoU gains under the 1% split are around 6%, 8%, and 10%, respectively. The results strongly verify the effectiveness of our framework and further highlight the importance of leveraging unlabeled data for LiDAR semantic segmentation.
Compare with State of the Arts. We compare LaserMix++ with recent arts from the literature, i.e., GPC [22], CRB [7], and Lim3D [11], and show results (using the Cylinder3D backbone) in Tab. II. As can be seen, LaserMix++ exhibits much better results than GPC [22] and [7] across various scenarios. The most recent work, Lim3D [11], achieves the best performance on ScribbleKITTI [7]. However, its results were obtained by voting from multiple augmented test-time ensembles, which created a different evaluation protocol from ours. In addition to the above, we also reproduced several popular algorithms [18, 67, 21] from the semi-supervised image segmentation domain. The results in Tab. II verify that these methods, albeit competitive in 2D tasks, only yield sub-par performance in the semi-supervised LiDAR semantic segmentation benchmark. This highlights the importance of exploiting the LiDAR data structure for data-efficient 3D scene understanding.
Compare with Fully Supervised Methods. As shown in Figure 1(c), the comparisons of LaserMix++ over the prevailing LiDAR semantic segmentation methods [33, 37, 35, 34] validate that our approaches are competitive to the fully supervised counterparts while only requiring 2$\times$ to 5$\times$ fewer annotations. Additionally, the results in Tab. II verify the strong augmentation and regularization ability of LaserMix++ again. Our approaches have yielded better results in the high-data regime and extreme low-data regime (i.e., $0.8\%$ human annotations on ScribbleKITTI [7]).
Qualitative Assessments. Figure 5 displays visualizations of the 3D scene segmentation results for different semi-supervised learning algorithms on the validation set of nuScenes [38], where each example covers a 50$\times$50 $\text{m}^{2}$ driving scene centered by the ego-vehicle. We observe that previous methods can only improve predictions in limited regions, while our approaches holistically eliminate false predictions in almost every region around the ego-vehicle. The consistency enlightened by our methods has yielded better 3D segmentation accuracy under annotation scarcity.
Enhancing Representation Learning Effects. Recent explorations on self-supervised LiDAR semantic segmentation exhibit promising representation learning effects during model pretraining [103, 104, 105, 76, 106, 77]. Such methods establish suitable self-supervised pretraining objectives and probe the qualitative of representation learning via few-shot fine-tuning. The results shown in Tab. IV verify that our approaches are effective in enhancing the effects of representation learning during fine-tuning. combined with Seal [77], LaserMix++ also achieves superior performance under full supervision (from 74.7% mIoU to 77.9% mIoU). This study further proves the versatility of our framework in handling different LiDAR semantic segmentation tasks.
Enhancing Out-of-Distribution Robustness. The ability to tackle scenarios that are not observed during the training time is crucial for a 3D scene understanding model, especially under autonomous driving context [6, 32]. The fine-grained manipulations of LiDAR point clouds in LaserMix and LaserMix++ have the potential to enrich the training distribution and, as a return, achieve better robustness. In this work, we conduct experiments on the Robo3D benchmark [31] to validate the robustness enhancement effect of our framework. As shown in Tab. III, our approaches consistently yield lower mCE scores for both LiDAR semantic segmentation and 3D object detection compared to other mixing-based techniques, such as Mix3D [101] and PolarMix [102].
Generalize to Image Segmentation. The proposed scene prior-based consistency regularization can be extended to image domains since they also reflect real-world driving scenarios. To validate this, we conduct experiments on Cityscapes [94] and show the results in Tab. V. As can be seen, our approaches achieved improved performance over the strong image segmentation baselines [18, 21, 19, 20]. Such a generalizability ensures our approaches are universally applicable to different sensor modalities.
(a) Mixing Strategy
(b) EMA
(c) Threshold
Figure 6: Ablation study on (a) Different mixing-based techniques used in point partition & mixing. (b) Different EMA decay rates between the Teacher and Student networks. (c) Different confidence thresholds $T$ used in the pseudo-label generation.
5.4 Ablation Study
Component Analyses. The ablation results in Tab. VI validate that each of the proposed components contributes significantly to the overall improvement in data-efficient 3D scene understanding. The mixing-based consistency regularization establishes a stable yet effective baseline across different datasets (Sec. 3.2). Meanwhile, using the Teacher network instead of the Student network to generate pseudo-labels tends to yield better results, as the formal temporally ensembles and encourages spatial consistency (Sec. 3.3). Moreover, integrating multi-modal interactions between LiDAR and cameras consistently improves performance, as encouraged by the camera-to-LiDAR feature distillation (Sec. 4.3) and the language-drive knowledge guidance (Sec. 4.4). It is worth noting that all model configurations have achieved superior results than the baseline MeanTeacher [18], which further emphasizes the effectiveness of our framework in tackling LiDAR segmentation under data-efficient learning setups.
Mixing Strategies. Another ablation experiment, depicted in Figure 6(a), compare the performance of LaserMix and LaserMix++ against traditional mixing techniques, i.e., MixUp [87], CutOut [111], CutMix [88], and Mix3D [101]. While MixUp and CutMix manipulate random points and areas respectively, they do not inherently leverage the structured nature of LiDAR data, where spatial relations significantly influence segmentation accuracy. CutMix shows improvement by utilizing the structural priors in scene segmentation; however, LaserMix and LaserMix++, which considers both area structure and precise spatial positioning, provides a more substantial boost in performance, outperforming CutMix by up to 3.3% mIoU. CutOut can be considered as setting $X_{\mathrm{out}}$ to a dummy filling instead of sampling from datasets, and it leads to a considerable performance drop from CutMix.
Alternative Heuristic Data Mixing. Exploring further, we evaluate different heuristic approaches to LiDAR scan partitioning, including azimuth and radius-based splits. As shown in Tab. VII, azimuth splits do not correlate well with semantic distributions, offering no performance gains. In contrast, finer granularity in mixing generally improves results until it reaches a threshold beyond which it disrupts semantic coherence. This observation supports our use of inclination-based partitioning in LaserMix, which maintains semantic integrity better than purely radial or azimuthal approaches.
Data Splitting Strategies. For a semi-supervised learning problem, the way of partitioning labeled and unlabeled data plays a crucial. We compare four different splitting strategies (the first three are shown in Figure 4) that meet the requirements of driving data collection and perception, i.e., random, uniform, sequential sampling strategies, and the recent visual-encouraged ST-RFD [11] strategy. As shown in Tab. VIII, methods under the random or uniform sampling achieved higher results than sequential sampling. This is because the former two introduce a more diverse sample distribution than sampling sequentially. ST-RFD which picks scans using additional image cues provides an even richer labeled data set. Under all four strategies, LaserMix++ consistently achieved better performance than the baseline LaserMix [10] and MeanTeacher [18] frameworks and is competitive with Lim3D [11] on ScribbleKITTI [7].
@@ 5.4 Ablation Study
Extension to Fully-Supervised Learning. The core principles of LaserMix, which involve partitioning and mixing LiDAR data to create diverse and challenging training samples, are not inherently limited to semi-supervised learning. These principles can be directly applied to fully-supervised settings, where the availability of labeled data allows for a more comprehensive exploration of the method’s potential. As shown in Tab. IX, our approach consistently improves by approximately 0.4% to 2.4% mIoU scores across different backbones. These results confirm that the core principles of our mixing-based methods, originally designed for semi-supervised tasks, can also significantly enhance performance in fully-supervised learning scenarios.

### v3

@@ 5.4 Ablation Study
Component Analyses. The ablation results in Table 5 validate that each of the proposed components contributes significantly to the overall improvement in data-efficient 3D scene understanding. The mixing-based consistency regularization establishes a stable yet effective baseline across different datasets (Section 3.2). Meanwhile, using the Teacher network instead of the Student network to generate pseudo-labels tends to yield better results, as the formal temporally ensembles and encourages spatial consistency (Section 3.3). Moreover, integrating multi-modal interactions between LiDAR and cameras consistently improves performance, as encouraged by the camera-to-LiDAR feature distillation (Section 4.3) and the language-driven knowledge guidance (Section 4.4). It is worth noting that all model configurations have achieved superior results than the baseline MeanTeacher [82], which further emphasizes the effectiveness of our framework in tackling LiDAR segmentation under data-efficient learning setups.
Mixing Strategies. Another ablation experiment, depicted in Figure 6(a), compares the performance of LaserMix and LaserMix++ against traditional mixing techniques, i.e., MixUp [100], CutOut [17], CutMix [97], and Mix3D [63]. While MixUp and CutMix manipulate random points and areas, respectively, they do not inherently leverage the structured nature of LiDAR data, where spatial relations significantly influence segmentation accuracy. CutMix shows improvement by utilizing the structural priors in scene segmentation; however, LaserMix and LaserMix++, which consider both area structure and precise spatial positioning, provide a more substantial boost in performance, outperforming CutMix by up to 3.3% mIoU. CutOut can be considered as setting $X_{\mathrm{out}}$ to a dummy filling instead of sampling from datasets, and it leads to a considerable performance drop from CutMix.
Alternative Heuristic Data Mixing. Exploring further, we evaluate different heuristic approaches to LiDAR scan partitioning, including azimuth and radius-based splits. As shown in Table 7, azimuth splits do not correlate well with semantic distributions, offering no performance gains. In contrast, finer granularity in mixing generally improves results until it reaches a threshold beyond which it disrupts semantic coherence. This observation supports our use of inclination-based partitioning in LaserMix, which maintains semantic integrity better than purely radial or azimuthal approaches.
(a) Mixing Strategy
(b) EMA
(c) Threshold
Figure 6: Ablation study on (a) Different mixing-based techniques used in point partition & mixing. (b) Different EMA decay rates between the Teacher and Student networks. (c) Different confidence thresholds $T$ that are used in the pseudo-label generation process.
Data Splitting Strategies. For a semi-supervised learning problem, the way of partitioning labeled and unlabeled data plays a crucial role. We compare four different splitting strategies (the first three are shown in Figure 4) that meet the requirements of driving data collection and perception, i.e., random, uniform, sequential sampling strategies, and the recent visual-encouraged ST-RFD [44] strategy. As shown in Table 8, methods under the random or uniform sampling achieved higher results than sequential sampling. This is because the former two introduce a more diverse sample distribution than sampling sequentially. ST-RFD, which picks scans using additional image cues, provides an even richer labeled data set. Under all four strategies, LaserMix++ consistently achieved better performance than the baseline LaserMix [38] and MeanTeacher [82] frameworks and is competitive with Lim3D [44] on ScribbleKITTI [83].
Table 9: Fully-supervised LiDAR semantic segmentation results on the val sets of SemanticKITTI [2], nuScenes [18], and ScribbleKITTI [83]. All IoU scores are given in percentage ($\%$). Symbol denotes that the model has been trained with LaserMix as the data augmentation during the training.
@@
Extension to Fully-Supervised Learning. The core principles of LaserMix, which involve partitioning and mixing LiDAR data to create diverse and challenging training samples, are not inherently limited to semi-supervised learning. These principles can be directly applied to fully-supervised settings, where the availability of labeled data allows for a more comprehensive exploration of the method’s potential. As shown in Table 9, our approach consistently improves by approximately 0.4% to 2.4% mIoU scores across different backbones. These results confirm that the core principles of our mixing-based methods, originally designed for semi-supervised tasks, can also significantly enhance performance in fully-supervised learning scenarios.